# Bridging Classical and Quantum Machine Learning: Knowledge Transfer from Classical to Quantum Neural Networks Using Knowledge Distillation

### v1

@@ 3.2 Quantum Superposition and the Mathematical Intuition Behind QNNs
A QNN can be represented by a parameterized quantum circuit (PQC) $U(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ denotes the vector of parameters. The output of the QNN for an input state $|\psi(\mathbf{x})\rangle$ prepared by a quantum feature map is given in  equation 8 by:
@@
The postulates of quantum mechanics [53] allow us to describe the state of a quantum system using a wave function or state vector in a Hilbert space. A quantum system in a state $|\psi\rangle$ can exist in a superposition of different basis states $\{|i\rangle\}$. This superposition is mathematically represented in  equation 9 as:
@@
The expectation value $\langle O\rangle$ of an observable $O$, which corresponds to a physical quantity that can be measured, is computed according to  equation 10 as:
@@
For an observable $O$ with eigenstates $|o_{k}\rangle$ and eigenvalues $o_{k}$, the expectation value takes the form of  equation 11:
@@
3.3 Quantum Gradient Descent and Parameter Optimization
In the training of quantum neural networks (QNNs), the optimization of parameters $\boldsymbol{\theta}$ is crucial. Due to the nature of quantum computing, traditional backpropagation as used in classical neural networks is not directly applicable. Instead, we utilize the parameter shift rule to compute gradients of quantum circuits.
@@ 3.3.1 Parameter Shift Rule
The parameter shift rule is a method to compute the gradient of an expectation value of an observable with respect to the parameters of a quantum circuit [54]. For a parameter $\theta_{j}$ and an observable $O$, the gradient can be calculated using the following  equation 12 as:
@@ 3.3.2 Quantum Gradient Descent
Once the gradients are computed, the parameters can be updated using a gradient descent algorithm. The update rule at each iteration $t$ is given by  equation 13:
@@ 3.3.3 Quantum Circuit Learning
The training process involves iteratively adjusting the parameters $\boldsymbol{\theta}$ to minimize the loss function $\mathcal{O}_{total}$. This is referred to as quantum circuit learning and is given by  equation 7. The optimization process leverages the quantum superposition and entanglement to explore the parameter space more efficiently than classical methods. The inherent probabilistic nature of quantum measurements is accounted for in the optimization loop, making the training resilient to the stochastic nature of quantum mechanics. By integrating these quantum-specific optimization techniques, we aim to exploit the full potential of quantum neural networks while addressing the challenges posed by the hardware limitations and the unique aspects of quantum computation.
3.4 Amplitude Encoding in Quantum Neural Networks
Amplitude encoding [59] is a technique used in quantum computing to map classical data into quantum states. This method leverages the ability of quantum states to exist in superpositions, allowing an exponential compression of data compared to classical representations.
A classical vector $\mathbf{x}\in\mathbb{R}^{n}$ can be normalized and encoded into the amplitudes of a quantum state $|\psi\rangle$ in a Hilbert space of $\log_{2}(n)$ qubits, given according to  equation 14:
$|\psi\rangle=\sum_{i=0}^{n-1}x_{i}|i\rangle,$ (14)
where $\{x_{i}\}$ are the normalized values of the classical vector $\mathbf{x}$, and $\{|i\rangle\}$ represents the computational basis states.
The normalization ensures that $\sum_{i=0}^{n-1}|x_{i}|^{2}=1$, which is a requirement for any valid quantum state.
Amplitude encoding is necessary for efficiently utilizing the limited number of qubits available in current quantum hardware. It allows the encoding of a classical dataset with $2^{n}$ features into a quantum system with only $n$ qubits, thus exponentially reducing the dimensionality of the data representation.
In practice, amplitude encoding can be implemented using various quantum gates and operations. For instance, the non-linear transformation given by $x^{\prime}=\pi\cdot\tanh(x)$ can be applied to preprocess the data before encoding it into the amplitudes. Here, $x^{\prime}$ denotes the transformed feature that is to be encoded into the quantum state. The hyperbolic tangent function ensures that the features are scaled to the range $[-\pi,\pi]$, which is suitable for encoding into quantum state amplitudes. Then, we can map these preprocessed classical features into the amplitudes of a quantum state according to  equation 15:
$|\psi(\mathbf{x})\rangle=\frac{1}{\|\mathbf{x}^{\prime}\|}\sum_{i=0}^{n-1}x^{\prime}_{i}|i\rangle,$ (15)
where $\|\mathbf{x}^{\prime}\|$ is the L2 norm of the vector $\mathbf{x}^{\prime}$.
The encoding process translates into the preparation of quantum states through a sequence of quantum gates that transform the initial state $|0\rangle^{\otimes n}$ into the desired superposition state $|\psi(\mathbf{x})\rangle$. The specific sequence of gates depends on the quantum hardware and the form of the feature vector. In quantum neural networks, amplitude encoding serves as the crucial first step that translates classical information into a form that quantum layers can process, leading to the exploitation of quantum properties such as superposition and entanglement in subsequent computations. By using amplitude encoding, we ensure that classical data is compatible with quantum processing, allowing the QNN to leverage the advantages of quantum computation in solving machine learning tasks.
4 Experiments and Results
4.1 Datasets Used
@@
In our experiments, we have employed two well-known datasets: MNIST and FashionMNIST. The datasets are shown in  figure 2. Two datasets are used to demonstrate the generalizability and broader application of the methodology in this study.
4.1.1 MNIST
@@
The MNIST dataset [13] is a large database of handwritten digits that is commonly used for training various image processing systems. It contains 60,000 training images and 10,000 testing images, each of which is a grayscale image of size 28$\times$28 pixels. The dataset is designed to allow researchers to benchmark their algorithms in classifying and recognizing handwritten numerical digits from 0 to 9. Due to its simplicity and the fact that it has been extensively studied, it serves as a standard for evaluating the performance of various machine learning techniques.
4.1.2 FashionMNIST
@@
The FashionMNIST dataset [14] is a more recent dataset comprising 60,000 training samples and 10,000 test samples. Each sample is a 28$\times$28 grayscale image, associated with a label from 10 classes. The classes represent different types of clothing and fashion items, such as T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. This dataset is intended to serve as a direct drop-in replacement for the original MNIST dataset for benchmarking machine learning algorithms, with the added complexity of distinguishing between different types of clothing.
Both datasets are preprocessed in a similar manner, normalized, and used as the basis for training and testing our models. By utilizing these datasets, we aim to show that our approach is not limited to a specific type of data or domain but can be applied to various image recognition tasks, highlighting its flexibility and wide applicability.
Figure 2: The datasets used in the experiments. (a) The MNIST dataset of written digits having 10 classes. (b) The FashionMNIST dataset of clothing containing 10 classes.
4.2 Dataset Filtration
In line with previous research that has addressed similar challenges in quantum machine learning [60, 9, 50], we filtered the MNIST and FashionMNIST datasets to reduce simulation time and focus on specific aspects of classification tasks. We created two subsets: MNIST 0-5, containing only the classes labeled from 0 to 5, and FashionMNIST 0-7, including classes from 0 to 7. Furthermore, within each class, we have randomly selected 150 samples to form the filtered datasets.
The filtration of datasets serves several purposes in our study. Firstly, it allows us to evaluate our method’s effectiveness across a range of classes while keeping computational resources manageable. Quantum training can be inherently slow and resource-intensive due to the current limitations of quantum hardware and the complex nature of quantum computations. By reducing the number of classes and samples, we aim to demonstrate the feasibility of training quantum models without the need for extensive resources, which is crucial for the practical adoption of quantum machine learning.
Choosing a smaller subset of samples is motivated by the need for computational efficiency. Quantum computers and simulators, particularly those accessible for research purposes, are subject to resource constraints such as limited qubits and short coherence times. These constraints make training on full datasets impractical. The selection of 150 samples per class allows for a diverse enough representation to train on and validate our models, ensuring that the performance metrics reflect the models’ ability to generalize from a limited sample size.
4.3 Classical Convolutional Neural Networks
@@
Convolutional Neural Networks (CNNs) have been paramount in the field of image recognition and classification. For the MNIST and FashionMNIST datasets, two architectures stand out due to their historical significance and performance: LeNet [61] and AlexNet [62]. In our experiments, we have used these two models individually and their ensemble by averaging the outputs of the final layers as classical teacher models. Both LeNet and AlexNet were originally intended for higher dimensional images. To accommodate the 28$\times$28 pixel grayscale images of MNIST and FashionMNIST, we have modified the input layers of these models to accept single-channel 28$\times$28 pixel input instead of their original input dimensions.
4.3.1 LeNet
@@
LeNet is one of the earliest convolutional neural networks that significantly impacted the field of deep learning. It is a much simpler network compared to AlexNet, consisting of two convolutional layers followed by two fully connected layers. LeNet is particularly well-suited for handwritten digit recognition tasks like MNIST due to its architecture being designed for this purpose. In our experiments, we have used the original LeNet architecture by changing the input layer to accept images of dimension 28$\times$28 instead of its original input image dimension of 32$\times$32, as shown in  figure 3. The intermediate layers were constructed according to the original architecture with no change.
@@
Figure 3: The modified LeNet architecture used in the implementation of the experiments. Two convolution layers, each followed by tanh activation and average pool layers were used. Finally, three fully connected layers were used to classify the images and get output probabilities from the model.
4.3.2 AlexNet
AlexNet is a deep CNN that marked a breakthrough in the ImageNet competition. It consists of five convolutional layers followed by three fully connected layers. AlexNet takes an RGB image of size 256$\times$256 as input and uses 11$\times$11 filters with a stride of 4 in the first convolutional layer. Since MNIST and FashionMNIST images are grayscale images of size 28$\times$28, the architecture of AlexNet was modified in the intermediate layers. figure 4 shows the followed architecture of AlexNet that was used instead: The modified architecture features an initial convolutional layer with 64 filters of size 3$\times$3, stride 1, and padding 1, followed by ReLU activation and a max pooling layer with a 2$\times$2 kernel and stride 2. The second layer has 192 filters of size 3$\times$3 with padding 1, followed by ReLU activation and another 2$\times$2 max pooling. The third convolutional layer consists of 384 filters with a kernel size of 3$\times$3 and padding of 1, followed by ReLU activation. This is followed by a series of fully connected layers: the first with 4096 neurons, the second also with 4096 neurons, and the final layer corresponding to the number of classes. Dropout is applied before the first and second fully connected layers to prevent overfitting. The model is adapted for single-channel grayscale images and the smaller spatial dimensions of the MNIST and FashionMNIST datasets. The use of ReLU activation functions, dropout, and overlapping pooling makes AlexNet robust against overfitting and capable of learning complex patterns in image data. In the context of MNIST and FashionMNIST, AlexNet’s architecture, although more complex than necessary for such simple datasets, provides an excellent baseline for performance due to its depth and capacity to learn intricate features.
@@ 4.3.2 AlexNet
Figure 4: The modified AlexNet architecture used in the implementation of the experiments. Three convolution layers, first two followed by ReLU activation and max pool layers were used. Finally, three fully connected layers were used to classify the images and get output probabilities from the model.
4.3.3 Ensemble of AlexNet and LeNet
The ensemble method combines the strengths of both LeNet and AlexNet to achieve better generalization. By averaging the output probabilities of the final layer from both models, the ensemble captures a more robust representation of the data. The ensembling can be mathematically represented by  equation 16 as:
$p_{ensemble}(y|x)=\frac{1}{2}\left(p_{LeNet}(y|x)+p_{AlexNet}(y|x)\right),$ (16)
where $p_{ensemble}(y|x)$ is the probability of the label $y$ given input $x$ for the ensemble, and $p_{AlexNet}(y|x)$ and $p_{LeNet}(y|x)$ are the probabilities given by the individual AlexNet and LeNet models, respectively.
This averaging process smooths out the predictions, making the ensemble less likely to overfit to noise in the dataset and often resulting in improved classification accuracy on test data. The use of these models and their ensemble as teacher models in our experiments aims to provide comprehensive learning signals for the student QNN, leveraging the classical CNNs’ ability to extract hierarchical features from image data.
4.4 Parameterized Quantum Circuits
In our study, we employed two Parameterized Quantum Circuits (PQCs): one with 3 qubits and another with 4 qubits. PQCs are central to quantum machine learning and variational algorithms, where the parameters (often denoted as $\theta$) of quantum gates are tuned to minimize a cost function. The PQCs that are constructed and used in our experiments are shown in  figure 5. The 3-qubit PQC is constructed using Hadamard gates (H gates) and rotation gates around the y-axis (Ry gates), which are parameterized by angles $\theta$. These angles are randomly initialized and optimized during the learning process. The circuit provides a compact yet expressive model that can capture the complexity needed for smaller-scale quantum tasks. Similarly, the 4-qubit PQC consists of a sequence of Ry gates with randomly initialized parameters $\theta$. With an additional qubit, the 4-qubit circuit can represent a larger state space, allowing it to capture more complex patterns and correlations in the data.
@@ 4.4 Parameterized Quantum Circuits
Figure 5: Parameterized Quantum Circuits used in the experiments. (a) A simple 3-qubit PQC with H and Ry gates parameterized by $\theta$. (b) A simple 4-qubit PQC with H and Ry gates parameterized by $\theta$.
The use of these two circuits is vital with respect to the flexibility and scalability of the methodology. A 3-qubit model demonstrates how our method performs with a minimal number of qubits, which is important for near-term quantum devices with limited qubit resources. The 4-qubit model, on the other hand, tests the capability of our method to scale to a slightly larger system, which is crucial for assessing performance in more demanding tasks. Together, they demonstrate generalizability and broader application potential across different quantum system sizes, which is essential for practical quantum machine learning implementations.
4.5 Performance Metrics

### v2

@@ 3.1 Quantum Superposition and the Mathematical Intuition Behind QNNs
A QNN can be represented by a parameterized quantum circuit (PQC) $U(\boldsymbol{\theta})$, where $\boldsymbol{\theta}$ denotes the vector of parameters. The output of the QNN for an input state $|\psi(\mathbf{x})\rangle$ prepared by a quantum feature map is given in Eq. 1 by:
@@
The postulates of quantum mechanics [56] allow us to describe the state of a quantum system using a wave function or state vector in a Hilbert space. A quantum system in a state $|\psi\rangle$ can exist in a superposition of different basis states $\{|i\rangle\}$. This superposition is mathematically represented in Eq. 2 as:
@@
The expectation value $\langle O\rangle$ of an observable $O$, which corresponds to a physical quantity that can be measured, is computed according to Eq. 3 as:
@@
For an observable $O$ with eigenstates $|o_{k}\rangle$ and eigenvalues $o_{k}$, the expectation value takes the form of Eq. 4:
@@
3.2 Data Encoding in QNNs
3.2.1 Angle Encoding
Angle encoding, also known as rotational encoding, is a method used in quantum computing to encode classical data into the parameters of quantum gates. This technique utilizes the rotation gates such as RX, RY, and RZ to map classical data points to angles of rotation, effectively embedding the data into the quantum state.
Given a classical vector $\mathbf{x}\in\mathbb{R}^{n}$, each element $x_{i}$ of the vector can be associated with a rotation angle for a corresponding quantum gate. The encoding can be represented as a sequence of operations on a quantum state $|\psi\rangle$ in a Hilbert space of $n$ qubits, as shown in Eq. 5:
$|\psi(\mathbf{x})\rangle=\bigotimes_{i=0}^{n-1}R_{\phi_{i}}(x_{i})|0\rangle,$ (5)
where $R_{\phi_{i}}$ denotes the rotation gate applied to the $i$-th qubit with $\phi_{i}$ being one of the axes $\{X,Y,Z\}$, and $x_{i}$ is the classical data point determining the rotation angle.
The angle encoding process involves applying a series of rotation gates to an initial quantum state, usually the zero state $|0\rangle^{\otimes n}$, transforming it into a state that represents the encoded classical data. The choice of rotation axis and the order of gate application can vary depending on the specific implementation and the desired properties of the encoded state.
Angle encoding is particularly useful in variational quantum circuits and quantum machine learning models, where the parameters of the quantum gates are optimized during the training process. This form of encoding allows the model to directly manipulate the quantum state using classical data, enabling the exploration of complex feature spaces through the quantum system’s inherent properties of superposition and entanglement.
$R_{\phi_{i}}(x_{i})=e^{-i\frac{x_{i}}{2}\sigma_{\phi_{i}}},$ (6)
where $\sigma_{\phi_{i}}$ is the Pauli operator corresponding to the rotation axis $\phi_{i}$, and $x_{i}$ is the angle of rotation derived from the classical data.
By employing angle encoding, a QNN can harness the full potential of quantum computation to process and learn from classical data, potentially leading to more powerful and efficient machine learning models.
3.2.2 Amplitude Encoding
Amplitude encoding is another technique used to embed classical data into the quantum state of a quantum system. Unlike angle encoding, which relies on rotation gates, amplitude encoding directly maps the classical data to the amplitudes of the quantum state’s basis vectors.
Given a classical vector $\mathbf{x}=(x_{0},x_{1},\ldots,x_{n-1})$, the amplitude encoding can be represented as:
$|\psi(\mathbf{x})\rangle=\sum_{i=0}^{2^{n}-1}x_{i}|i\rangle,$ (7)
where $|i\rangle$ represents the computational basis state corresponding to the binary representation of the index $i$, and the amplitudes $x_{i}$ are derived from the classical data $\mathbf{x}$.
The normalization condition for the quantum state must be satisfied:
$\sum_{i=0}^{2^{n}-1}|x_{i}|^{2}=1.$ (8)
Amplitude encoding allows for a direct mapping of classical data to the quantum state, potentially enabling more efficient processing and manipulation of the encoded information. However, it comes with the challenge of ensuring the normalization condition is met, which may require additional preprocessing or normalization steps.
3.2.3 Qubit Encoding
Qubit encoding, also known as basis encoding, represents classical data by associating each basis state of a quantum system with a specific data point. This encoding method is particularly suitable for discrete or categorical data.
Consider a classical dataset $\mathcal{D}=\{d_{1},d_{2},\ldots,d_{N}\}$ consisting of $N$ distinct data points. Each data point $d_{i}$ can be mapped to a computational basis state $|i\rangle$ of an $n$-qubit system, where $n=\lceil\log_{2}N\rceil$.
The encoding process can be represented as:
$\mathcal{E}:\mathcal{D}\rightarrow\{|0\rangle,|1\rangle,\ldots,|2^{n}-1\rangle\},$ (9)
where $\mathcal{E}$ is the encoding function that maps each data point $d_{i}$ to the corresponding computational basis state $|i\rangle$.
To encode a specific data point $d_{i}$, the quantum state is prepared in the basis state $|i\rangle$ using a sequence of quantum gates, such as Hadamard gates and controlled-NOT gates.
Qubit encoding is particularly useful when dealing with discrete or categorical data, as it allows for a direct mapping of data points to the computational basis states. However, for continuous data, additional preprocessing or discretization steps may be required before encoding.
3.3 QNN Training
In the training of QNNs, the optimization of parameters $\boldsymbol{\theta}$ is crucial. Due to the nature of quantum computing, traditional backpropagation as used in classical neural networks is not directly applicable. Instead, we utilize the parameter shift rule to compute gradients of quantum circuits.
@@ 3.3.1 Parameter Shift Rule
The parameter shift rule is a method to compute the gradient of an expectation value of an observable with respect to the parameters of a quantum circuit [57]. For a parameter $\theta_{j}$ and an observable $O$, the gradient can be calculated using the following Eq. 10 as:
@@ 3.3.2 Quantum Gradient Descent
Once the gradients are computed, the parameters can be updated using a gradient descent algorithm. The update rule at each iteration $t$ is given by Eq. 11:
@@ 3.3.3 Quantum Circuit Learning
The training process involves iteratively adjusting the parameters $\boldsymbol{\theta}$ to minimize the loss function $\mathcal{O}_{total}$. This is referred to as quantum circuit learning and is given by Eq. 24. The optimization process leverages the quantum superposition and entanglement to explore the parameter space more efficiently than classical methods. The inherent probabilistic nature of quantum measurements is accounted for in the optimization loop, making the training resilient to the stochastic nature of quantum mechanics. By integrating these quantum-specific optimization techniques, we aim to exploit the full potential of QNNs while addressing the challenges posed by the hardware limitations and the unique aspects of quantum computation.
3.3.4 Quantum Measurement and Quantum to Classical Decoding
After the optimization of the QNN’s parameters $\boldsymbol{\theta}$, the final step is to measure the quantum state and decode the quantum output into a classical form suitable for the target task, such as classification or regression. According to the principles of quantum mechanics, the measurement of a quantum system causes the system to collapse into one of the possible eigenstates of the measured observable. The probability of obtaining a particular eigenvalue is determined by the squared modulus of the corresponding eigenstate’s amplitude in the quantum state before measurement.
In QNNs, the observable $O$ is chosen such that its eigenvalues correspond to the desired classical outputs. For example, in a binary classification task, the observable can be defined with eigenvalues $\{0,1\}$ representing the two classes. The expectation value $\langle O\rangle$ obtained from the QNN circuit, as given in Eq. 3, represents the weighted average of the observable’s eigenvalues, where the weights are the probabilities of obtaining each eigenvalue upon measurement.
However, due to the inherent probabilistic nature of quantum mechanics, a single measurement of the quantum state may not provide an accurate estimate of the expectation value. To overcome this limitation, multiple measurements, or “shots," are performed on identically prepared quantum states.
Let $N$ be the number of shots, and $n_{k}$ be the number of times the eigenvalue $o_{k}$ is obtained after $N$ measurements. The estimated probability $\hat{P}(o_{k})$ of obtaining the eigenvalue $o_{k}$ is given by:
$\hat{P}(o_{k})=\frac{n_{k}}{N}.$ (12)
The estimated expectation value $\langle\hat{O}\rangle$ can then be computed as:
$\langle\hat{O}\rangle=\sum_{k}o_{k}\hat{P}(o_{k})=\frac{1}{N}\sum_{i=1}^{N}o_{i},$ (13)
where $o_{i}$ is the eigenvalue obtained in the $i$-th measurement.
By performing a large number of shots $N$, the estimated expectation value $\langle\hat{O}\rangle$ converges to the true expectation value $\langle O\rangle$ according to the law of large numbers. The estimated probabilities $\hat{P}(o_{k})$ can then be interpreted as the QNN’s output for the corresponding classes or regression values, enabling the decoding of the quantum output into a classical form suitable for the target task.
4 Materials and Methods
4.1 Materials
4.1.1 Datasets
@@
In our experiments, we have employed three well-known datasets: MNIST, FashionMNIST, and CIFAR10. The datasets are shown in Fig. 2, Fig. 3, and Fig. 4. These datasets are used to demonstrate the generalizability and broader application of the methodology in this study.
@@
MNIST. The MNIST dataset [12] is a large database of handwritten digits that is commonly used for training various image processing systems. It contains 60,000 training images and 10,000 testing images, each of which is a grayscale image of size 28$\times$28 pixels. The dataset is designed to allow researchers to benchmark their algorithms in classifying and recognizing handwritten numerical digits from 0 to 9. Due to its simplicity and the fact that it has been extensively studied, it serves as a standard for evaluating the performance of various machine learning techniques.
Figure 2: The MNIST dataset of handwritten digits having 10 classes
Figure 3: The FashionMNIST dataset of clothing items with 10 classes
@@
FashionMNIST. The FashionMNIST dataset [13] is a more recent dataset comprising 60,000 training samples and 10,000 test samples. Each sample is a 28$\times$28 grayscale image, associated with a label from 10 classes. The classes represent different types of clothing and fashion items, such as T-shirts/tops, trousers, pullovers, dresses, coats, sandals, shirts, sneakers, bags, and ankle boots. This dataset is intended to serve as a direct drop-in replacement for the original MNIST dataset for benchmarking machine learning algorithms, with the added complexity of distinguishing between different types of clothing.
Figure 4: The CIFAR10 dataset with RGB image classes for image classification
CIFAR10. The CIFAR10 dataset [14] consists of 60,000 32$\times$32 color images in 10 different classes. The classes include airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks. There are 6,000 images per class, divided into 50,000 training images and 10,000 test images. The dataset is another benchmark in the field of machine learning and computer vision, providing a more challenging dataset compared to MNIST and FashionMNIST due to its color content and more complex image patterns.
4.1.2 Classical Convolutional Neural Networks
@@
Convolutional Neural Networks (CNNs) have been paramount in the field of image recognition and classification. For the MNIST and FashionMNIST datasets, two architectures stand out due to their historical significance and performance: LeNet [62] and AlexNet [63]. In our experiments, we have used these two models as classical teacher models for the grayscale images. Both LeNet and AlexNet were originally intended for higher dimensional images. To accommodate the 28$\times$28 pixel grayscale images of MNIST and FashionMNIST, we have modified the input layers of these models to accept single-channel 28$\times$28 pixel input instead of their original input dimensions.
For the CIFAR10 dataset, which consists of 32$\times$32 pixel color images, we have utilized pre-trained models of ResNet50 [64], DenseNet121 [65], and EfficientNetB0 [66] without any architectural modifications. These models are well-suited for handling the higher complexity and information present in the CIFAR10 images.
@@
Modified LeNet Architecture. LeNet is one of the earliest convolutional neural networks that significantly impacted the field of deep learning. It is a much simpler network compared to AlexNet, consisting of two convolutional layers followed by two fully connected layers. LeNet is particularly well-suited for handwritten digit recognition tasks like MNIST due to its architecture being designed for this purpose. In our experiments, we have used the original LeNet architecture by changing the input layer to accept images of dimension 28$\times$28 instead of its original input image dimension of 32$\times$32, as shown in Fig. 6. The intermediate layers were constructed according to the original architecture with no change.
@@
Figure 5: The modified LeNet architecture used in the implementation of the experiments. Two convolution layers, each followed by tanh activation and average pool layers were used. Finally, three fully connected layers were used to classify the images and get output probabilities from the model
@@
Figure 6: The modified AlexNet architecture used in the implementation of the experiments. Three convolution layers, first two followed by ReLU activation and max pool layers were used. Finally, three fully connected layers were used to classify the images and get output probabilities from the model
Modified AlexNet Architecture. AlexNet is a deep CNN that marked a breakthrough in the ImageNet competition. It consists of five convolutional layers followed by three fully connected layers. AlexNet takes an RGB image of size 256$\times$256 as input and uses 11$\times$11 filters with a stride of 4 in the first convolutional layer. Since MNIST and FashionMNIST images are grayscale images of size 28$\times$28, the architecture of AlexNet was modified in the intermediate layers. Fig. 6 shows the followed architecture of AlexNet that was used instead: The modified architecture features an initial convolutional layer with 64 filters of size 3$\times$3, stride 1, and padding 1, followed by ReLU activation and a max pooling layer with a 2$\times$2 kernel and stride 2. The second layer has 192 filters of size 3$\times$3 with padding 1, followed by ReLU activation and another 2$\times$2 max pooling. The third convolutional layer consists of 384 filters with a kernel size of 3$\times$3 and padding of 1, followed by ReLU activation. This is followed by a series of fully connected layers: the first with 4096 neurons, the second also with 4096 neurons, and the final layer corresponding to the number of classes. Dropout is applied before the first and second fully connected layers to prevent overfitting. The model is adapted for single-channel grayscale images and the smaller spatial dimensions of the MNIST and FashionMNIST datasets. The use of ReLU activation functions, dropout, and overlapping pooling makes AlexNet robust against overfitting and capable of learning complex patterns in image data. In the context of MNIST and FashionMNIST, AlexNet’s architecture, although more complex than necessary for such simple datasets, provides an excellent baseline for performance due to its depth and capacity to learn intricate features.
@@
Figure 7: The 4-qubit PQC with Rx, Ry and Rz gates parameterized by $\theta$ used as student model for image classification
Figure 8: The 8-qubit PQC with Rx, Ry and Rz gates parameterized by $\theta$ used as student model for image classification
4.1.3 Parameterized Quantum Circuits
In our study, we employed two Parameterized Quantum Circuits (PQCs): one with 4 qubits and another with 8 qubits. PQCs are central to QML and variational algorithms, where the parameters (often denoted as $\theta$) of quantum gates are tuned to minimize a cost function. The PQCs that are constructed and used in our experiments are shown in Fig. 8 and Fig. 8. The PQCs are constructed using rotation gates Rx, Ry and Rz, which are parameterized by angles $\theta$. These angles are randomly initialized and optimized during the learning process. The 4-qubit circuit provides a compact yet expressive model that can capture the complexity needed for smaller-scale quantum tasks. With four additional qubits, the 8-qubit circuit can represent a larger state space, allowing it to capture more complex patterns and correlations in the data.
4.1.4 Performance Metrics